# On the Curses of Future and History in Future-dependent Value Functions for Off-policy Evaluation

### v1

@@ Abstract
We study off-policy evaluation (OPE) in partially observable environments with complex observations, with the goal of developing estimators whose guarantee avoids exponential dependence on the horizon. While such estimators exist for MDPs and POMDPs can be converted to history-based MDPs, their estimation errors depend on the state-density ratio for MDPs which becomes history ratios after conversion, an exponential object. Recently, Uehara et al. (2022) proposed future-dependent value functions as a promising framework to address this issue, where the guarantee for memoryless policies depends on the density ratio over the latent state space. However, it also depends on the boundedness of the future-dependent value function and other related quantities, which we show could be exponential-in-length and thus erasing the advantage of the method. In this paper, we discover novel coverage assumptions tailored to the structure of POMDPs, such as outcome coverage and belief coverage. These assumptions not only enable polynomial bounds on the aforementioned quantities, but also lead to the discovery of new algorithms with complementary properties.
off-policy evaluation, reinforcement learning, POMDP
@@ 1 Introduction and Related Works
Off-policy evaluation (OPE) is the problem of estimating the return of a new evaluation policy $\pi_{e}$ based on historical data, which is typically collected using a different policy $\pi_{b}$ (the behavior policy). OPE plays a central role in the pipeline of offline reinforcement learning (RL), but is also notoriously difficult. Among the major approaches, importance sampling (IS) and its variants (Precup et al., 2000; Jiang & Li, 2016) provide unbiased and/or asymptotically correct estimation using the cumulative importance weights: given a trajectory of observations and actions
$o_{1},a_{1},o_{2},a_{2},\ldots,o_{H},a_{H},$
the cumulative importance weight is $\prod_{h=1}^{H}\frac{\pi_{e}(a_{h}|o_{h})}{\pi_{b}(a_{h}|o_{h})}$, whose variance grows exponentially with the horizon, unless $\pi_{b}$ and $\pi_{e}$ are very close in their action distributions. In the language of offline RL theory (Chen & Jiang, 2019; Xie & Jiang, 2021; Yin & Wang, 2021), the boundedness of the cumulative importance weights is the coverage assumption required by IS, a very stringent one.
@@
Alternatively, algorithms such as Fitted-Q Evaluation (FQE; Ernst et al., 2005; Munos & Szepesvári, 2008; Le et al., 2019) and Marginalized Importance Sampling (MIS; Liu et al., 2018; Xie et al., 2019; Nachum et al., 2019; Uehara et al., 2020) enjoy more favorable coverage assumptions, at the cost of function-approximation biases. Instead of requiring bounded cumulative importance weights, FQE and MIS only require that of the state-density ratios, which can be substantially smaller (Chen & Jiang, 2019; Xie & Jiang, 2021). That is, when the environment satisfies the Markov assumption (namely $o_{h}$ is a state), the guarantees of FQE and MIS only depend on the range of $d^{\pi_{e}}_{h}(o_{h})/d^{\pi_{b}}_{h}(o_{h})$, where $d^{\pi_{e}}_{h}$ and $d^{\pi_{b}}_{h}$ is the marginal distribution of $o_{h}$ under $\pi_{e}$ and $\pi_{b}$, respectively.
@@
In this paper, we study the non-Markov setting, which is ubiquitous in real-world applications. Such environments are typically modeled as Partially Observable Markov Decision Processes (POMDPs) (Kaelbling et al., 1998). Despite the more general formulation, one can reduce a POMDP to an MDP, making algorithms for MDPs applicable: we can simply define an equivalent MDP, with its state being the history of the original POMDP, $(o_{1},a_{1},\ldots,o_{h})$. Unfortunately, a close inspection reveals the problem: the state-density ratio after conversion is
$\frac{d^{\pi_{e}}_{h}(o_{1},a_{1},\ldots,o_{h})}{d^{\pi_{b}}_{h}(o_{1},a_{1},\ldots,o_{h})}=\prod_{h^{\prime}=1}^{h-1}\frac{\pi_{e}(a_{h^{\prime}}|o_{h^{\prime}})}{\pi_{b}(a_{h^{\prime}}|o_{h^{\prime}})},$
which is exactly the cumulative importance weights in IS and thus also an exponential object!
@@
To address this issue, Uehara et al. (2022) recently proposed a promising framework called future-dependent value functions (or FDVF for short). Notably, their coverage assumption is the boundedness of density ratios between $\pi_{e}$ and $\pi_{b}$ over the latent state for memoryless policies. This is as if we were dealing directly with the latent MDP underlying the POMDP, a perhaps best possible scenario. Nevertheless,
have we achieved exponential-free OPE in POMDPs?
@@
The answer to this question turns out to be nontrivial. In addition to the latent-state coverage parameter, the guarantee in Uehara et al. (2022) also depends on other quantities that are less interpretable. Among them, the boundedness of FDVF itself—a concept central to this framework—is unclear, and we show that a natural construction yields an upper bound that still scales with the cumulative importance weights, thus possibly erasing the superiority of the framework over IS or MDP-reduction.
@@
    With another novel concept called belief coverage (Section 5.1), we establish fully polynomial estimation guarantee for the algorithm in Uehara et al. (2022). The discovery of belief coverage also leads to a novel algorithm (Section 5.3) that is analogous to MIS for MDPs.
@@
In addition, certain obstacles we encounter are closely related to quantities in online RL for POMDPs (e.g., the $\alpha$-revealing parameter of Liu et al. (2022); Chen et al. (2023)), and our solution may shed light on these problems.
@@ 2 Preliminaries
We consider a finite-horizon POMDP $\left\langle{H,{\mathcal{S}}=\bigcup_{h=1}^{H}{\mathcal{S}}_{h},\mathcal{A},\mathcal{O}=\bigcup_{h=1}^{H}\mathcal{O}_{h},R,\mathbb{O},\mathbb{T},d_{1}}\right\rangle$, where $H$ is the horizon, ${\mathcal{S}}_{h}$ is the latent state space at step $h$ with $|{\mathcal{S}}_{h}|=S$, $\mathcal{A}$ is the action space with $|\mathcal{A}|=A$, $\mathcal{O}_{h}$ is the observation space at step $h$ with $|\mathcal{O}_{h}|=O$ , $R:\mathcal{O}\times\mathcal{A}\rightarrow[0,1]$ is the reward function, [^1] Some POMDP formulation assumes reward is a function of $s_{h}$, which can be absorbed into the observation since it is visible. $\mathbb{O}:{\mathcal{S}}\rightarrow\Delta(\mathcal{O})$ is the emission dynamics with $\mathbb{O}(\cdot|s_{h})$ supported on $\mathcal{O}_{h}$ for $s_{h}\in{\mathcal{S}}_{h}$, $\mathbb{T}:{\mathcal{S}}\times\mathcal{A}\to\Delta({\mathcal{S}})$ is the dynamics ($\mathbb{T}(\cdot|s_{h},a_{h})$ is supported on ${\mathcal{S}}_{h+1}$), and $d_{1}\in\Delta({\mathcal{S}}_{1})$ is the initial latent state distribution. For mathematical convenience we assume all the spaces are finite and discrete, but the cardinality of $\mathcal{O}_{h}$, $O$, can be arbitrarily large. A trajectory (or episode) is sampled as $s_{1}\sim d_{1}$, then $o_{h}\sim\mathbb{O}(\cdot|s_{h})$, $r_{h}=R(o_{h},a_{h})$, $s_{h+1}\sim\mathbb{T}(\cdot|s_{h},a_{h})$ for $1\leq h\leq H$, with $a_{1:H}$ decided by the decision-making agent, and the episode terminates after $a_{H}$. $s_{1:H}$ are latent and not observable to the agent.
@@
Given an episode $o_{1},a_{1},\ldots,o_{H},a_{H}$ and a time step $h$ of interest, it will be convenient to rewrite the episode as
$(\tau_{h}\,,\,\overbrace{o_{h}\,,\,a_{h}\,,\,f_{h+1}}^{f_{h}}).$
Here $\tau_{h}=(o_{1},a_{1},\ldots o_{h-1},a_{h-1})\in\mathcal{H}_{h}:=\prod_{h^{\prime}=1}^{h-1}(\mathcal{O}_{h}\times\mathcal{A})$ denotes the historical observation-action sequence (or simply history) prior to step $h$, and $f_{h+1}=(o_{h+1},a_{h+1},\ldots,o_{H},a_{H})\in\mathcal{F}_{h+1}:=\prod_{h^{\prime}=h+1}^{H}(\mathcal{O}_{h^{\prime}}\times\mathcal{A})$ denotes the future after step $h$. This format will be convenient for reasoning about the system dynamics at step $h$. We use $\mathcal{H}=\bigcup_{h=1}^{H}\mathcal{H}_{h}$ to denote the entire history domain and use $\mathcal{F}=\bigcup_{h=1}^{H}\mathcal{F}_{h}$ to denote the entire future domain. This way we can use stationary notation for functions over ${\mathcal{S}},\mathcal{O},\mathcal{H},\mathcal{F}$, where the time step can be identified from the function input (e.g., $R(o_{h},a_{h})$), and functions with time-step subscripts refer to their restriction to the $h$-th step input space, often treated as a vector (e.g., $R_{h}\in\mathbb{R}^{\mathcal{O}_{h}\times\mathcal{A}}$).
Memoryless Policy.
A policy $\pi:\bigcup_{h=1}^{H}(\mathcal{H}_{h}\times\mathcal{O}_{h})\rightarrow\Delta(\mathcal{A})$ specifies the action probability conditioned on the past observation-action sequence. For most of the paper, we will restrict ourselves to memory-less (or reactive) policies that only depends on the current observation $o_{h}$. For any $\pi$, we use $\Pr_{\pi}$ and $\mathbb{E}_{\pi}$ for the probabilities and expectations under episodes generated by $\pi$, and define $J(\pi)$ as the expected cumulative return: $J(\pi):=\mathbb{E}_{\pi}\left[\sum_{h=1}^{H}R(o_{h},a_{h})\right]$. For memoryless $\pi$, we also define $V_{{\mathcal{S}}}^{\pi}(s_{h})$ as the latent state value function at $s_{h}$: $V_{{\mathcal{S}}}^{\pi}(s_{h}):=\mathbb{E}_{\pi}\left[\sum_{h^{\prime}=h}^{H}R(o_{h^{\prime}},a_{h^{\prime}})\mid s_{h}\right]\in[0,H]$. We will use $d^{\pi}$ to denote the marginal distribution of variables induced by $\pi$, mostly for the latent state $s_{h}$.
@@
In OPE, the goal is to estimate $J(\pi_{e})$ using $n$ data trajectories $\mathcal{D}=\{(o_{1}^{(i)},a_{1}^{(i)},r_{1}^{(i)},\ldots,o_{H}^{(i)},a_{H}^{(i)},r_{H}^{(i)}):i\in[n]\}$ collected using $\pi_{b}$. We write $\mathbb{E}_{\mathcal{D}}[\cdot]$ to denote empirical approximation of expectation using $\mathcal{D}$.
Define the one-step action probability ratio $\mu(o_{h},a_{h}):=\frac{\pi_{e}(a_{h}\mid o_{h})}{\pi_{b}(a_{h}\mid o_{h})}$. We make the following assumption throughout:
@@
We assume $\mu(o_{h},a_{h})$ is known and $\max_{h}\max_{o_{h},a_{h}}\mu(o_{h},a_{h})\leq C_{\mu}$.
This is a standard assumption in the OPE literature, and is needed by IS and value-based estimators that model state value functions (Jiang & Li, 2016; Liu et al., 2018).
@@
For mathematical conveniences, we make the following assumptions throughout:
@@
$\forall h\in[H]$,
1. $\textrm{rank}(M_{\mathcal{H},h})=\textrm{rank}(M_{\mathcal{F},h})=S$.
2. $\forall f_{h}$,  $\Pr_{\pi_{b}}(f_{h})>0$; $\forall o_{h},a_{h}$, $R(o_{h},a_{h})>0$.
These assumptions are merely for simplifying technical presentations, which allow us to invert certain covariance matrices and avoid $/0$ 0 situations. When they do not hold, the corner cases can be handled with some extra care and do not affect our main results and implications, and similar assumptions are common in RL theory (Zhan et al., 2022).
@@ 3 Future-dependent Value Functions
In this section, we provide a recap of FDVFs and translate the main result of Uehara et al. (2022) into the finite-horizon setting, which is mathematically cleaner and more natural in many aspects; see Appendix A.2 for further discussion.
@@
The central idea in Uehara et al. (2022) is to find observable proxies of $V_{{\mathcal{S}}}^{\pi_{e}}(s_{h})$, which takes future as inputs:
Definition 3 (Future-dependent value functions).
A future-dependent value function $V_{\mathcal{F}}:\mathcal{F}\to\mathbb{R}$, where $\mathcal{F}:=\bigcup_{h}\mathcal{F}_{h}$, is any function that satisfies the following: $\forall s_{h}$,
$\mathbb{E}_{\pi_{b}}\left[V_{\mathcal{F}}(f_{h})\mid s_{h}\right]=V_{{\mathcal{S}}}^{\pi_{e}}(s_{h}).$
Equivalently, in matrix form, we have $\forall h$, $M_{\mathcal{F},h}\times V_{\mathcal{F},h}=V_{{\mathcal{S}},h}^{\pi_{e}}$. Recall our convention, that $V_{\mathcal{F},h}\in\mathbb{R}^{|\mathcal{F}_{h}|}$ is $V_{\mathcal{F}}$ restricted to $\mathcal{F}_{h}$, and $V_{{\mathcal{S}},h}^{\pi_{e}}$ is defined similarly.
A few important properties of FDVFs:
  1. 1.
    A FDVF $V_{\mathcal{F}}$ is a property of $\pi_{e}$, but also depends on $\pi_{b}$. As we will see later in Section 4, the boundedness of $V_{\mathcal{F}}$ will depend on certain notion of coverage of $\pi_{b}$ over $\pi_{e}$.
  2. 2.
    Given $\pi_{b}$ and $\pi_{e}$, the FDVF $V_{\mathcal{F}}$ is generally not unique. In fact, Eq.(6) is generally an underdetermined linear system ($S\ll|\mathcal{F}_{h}|$), which can yield many solutions. As we see below, it suffices to model any one of the solutions. Thus, from now on, when we talk about the boundedness of $V_{\mathcal{F}}$, we always consider the $V_{\mathcal{F}}$ with the smallest range among all solutions to Eq.(6).
@@
Given any $V:\mathcal{F}\to\mathbb{R}$, the Bellman residual on state $s_{h}$ is defined as: [^2] We let $V(f_{H+1})\equiv 0$ for all $V$ to be considered, so that we do not need to handle the $H$-th step separately in the definitions.
@@
$(\mathcal{B}^{\mathcal{H}}V)(\tau_{h}):=$ $~{}\mathbb{E}_{\begin{subarray}{c}a_{h}\sim\pi_{e}\\ a_{h+1:H}\sim\pi_{b}\end{subarray}}[r_{h}+V(f_{h+1})\mid\tau_{h}]-\mathbb{E}_{\pi_{b}}[V(f_{h})\mid\tau_{h}]$ (3)
$=$ $~{}\langle\mathbf{b}(\tau_{h}),\mathcal{B}^{{\mathcal{S}}}_{h}V\rangle,$ (4)
where $\mathcal{B}^{{\mathcal{S}}}_{h}V\in\mathbb{R}^{S}$ is $\mathcal{B}^{{\mathcal{S}}}V$ restricted to ${\mathcal{S}}_{h}$.
@@
For any $\pi_{e}$, $\pi_{b}$, and $V:\mathcal{F}\to\mathbb{R}$,
$J(\pi_{e})-\mathbb{E}_{\pi_{b}}[V(f_{1})]$ $=\textstyle\sum_{h=1}^{H}\mathbb{E}_{\pi_{e}}\left[(\mathcal{B}^{{\mathcal{S}}}V)(s_{h})\right].$
See proof in Appendix B.1. As the lemma shows, any $V$ with small $\mathcal{B}^{{\mathcal{S}}}V$ (such as $V_{\mathcal{F}}$, since $\mathcal{B}^{{\mathcal{S}}}V_{\mathcal{F}}\equiv 0$) can be used to estimate $J(\pi_{e})$ via $\mathbb{E}_{\pi_{b}}[V(f_{1})]$. But again, $\mathcal{B}^{{\mathcal{S}}}$ operates on ${\mathcal{S}}$ which is unobserved, and we turn to its proxy $\mathcal{B}^{\mathcal{H}}$, which are linear measures of $\mathcal{B}^{{\mathcal{S}}}$ (Eq.3). More concretely, Uehara et al. (2022) proposed to estimate $\mathbb{E}_{\pi_{b}}[(\mathcal{B}^{\mathcal{H}}V)^{2}]$ using an additional helper class $\Xi:\mathcal{H}\to\mathbb{R}$ to handle the double-sampling issue (Antos et al., 2008; Dai et al., 2018):

### v2

# On the Curses of Future and History in Future-dependent Value Functions for OPE

@@ Abstract
We study off-policy evaluation (OPE) in partially observable environments with complex observations, with the goal of developing estimators whose guarantee avoids exponential dependence on the horizon. While such estimators exist for MDPs and POMDPs can be converted to history-based MDPs, their estimation errors depend on the state-density ratio for MDPs which becomes history ratios after conversion, an exponential object. Recently, Uehara et al. (2022a) proposed future-dependent value functions as a promising framework to address this issue, where the guarantee for memoryless policies depends on the density ratio over the latent state space. However, it also depends on the boundedness of the future-dependent value function and other related quantities, which we show could be exponential-in-length and thus erasing the advantage of the method. In this paper, we discover novel coverage assumptions tailored to the structure of POMDPs, such as outcome coverage and belief coverage, which enable polynomial bounds on the aforementioned quantities. As a side product, our analyses also lead to the discovery of new algorithms with complementary properties.
@@ 1 Introduction and Related Works
Off-policy evaluation (OPE) is the problem of estimating the return of a new evaluation policy $\pi_{e}$ based on historical data, which is typically collected using a different policy $\pi_{b}$ (the behavior policy). OPE plays a central role in the pipeline of offline reinforcement learning (RL), but is also notoriously difficult. Among the major approaches, importance sampling (IS) and its variants Precup et al. (2000); Jiang and Li (2016) provide unbiased and/or asymptotically correct estimation using the cumulative importance weights: given a trajectory of observations and actions $o_{1},a_{1},o_{2},a_{2},\ldots,o_{H},a_{H},$ the cumulative importance weight is $\prod_{h=1}^{H}\frac{\pi_{e}(a_{h}|o_{h})}{\pi_{b}(a_{h}|o_{h})}$, whose variance grows exponentially with the horizon, unless $\pi_{b}$ and $\pi_{e}$ are very close in their action distributions. In the language of offline RL theory (Chen and Jiang, 2019; Xie and Jiang, 2021; Yin and Wang, 2021), the boundedness of the cumulative importance weights is the coverage assumption required by IS, a very stringent one.
@@
Alternatively, algorithms such as Fitted-Q Evaluation (FQE; Ernst et al., 2005; Munos and Szepesvári, 2008; Le et al., 2019) and Marginalized Importance Sampling (MIS; Liu et al., 2018; Xie et al., 2019; Nachum et al., 2019; Uehara et al., 2020) enjoy more favorable coverage assumptions, at the cost of function-approximation biases. Instead of requiring bounded cumulative importance weights, FQE and MIS only require that of the state-density ratios, which can be substantially smaller (Chen and Jiang, 2019; Xie and Jiang, 2021). That is, when the environment satisfies the Markov assumption (namely $o_{h}$ is a state), the guarantees of FQE and MIS only depend on the range of $d^{\pi_{e}}_{h}(o_{h})/d^{\pi_{b}}_{h}(o_{h})$, where $d^{\pi_{e}}_{h}$ and $d^{\pi_{b}}_{h}$ is the marginal distribution of $o_{h}$ under $\pi_{e}$ and $\pi_{b}$, respectively.
@@
In this paper, we study the non-Markov setting, which is ubiquitous in real-world applications. Such environments are typically modeled as Partially Observable Markov Decision Processes (POMDPs) Kaelbling et al. (1998). Despite the more general formulation, one can reduce a POMDP to an MDP, making algorithms for MDPs applicable: we can simply define an equivalent MDP, with its state being the history of the original POMDP, $(o_{1},a_{1},\ldots,o_{h})$. Unfortunately, a close inspection reveals the problem: the state-density ratio after conversion is $\frac{d^{\pi_{e}}_{h}(o_{1},a_{1},\ldots,o_{h})}{d^{\pi_{b}}_{h}(o_{1},a_{1},\ldots,o_{h})}=\prod_{h^{\prime}=1}^{h-1}\frac{\pi_{e}(a_{h^{\prime}}|o_{h^{\prime}})}{\pi_{b}(a_{h^{\prime}}|o_{h^{\prime}})},$ which is exactly the cumulative importance weights in IS and thus also an exponential object!
@@
To address this issue, Uehara et al. (2022a) recently proposed a promising framework called future-dependent value functions (or FDVF for short). Notably, their coverage assumption is the boundedness of density ratios between $\pi_{e}$ and $\pi_{b}$ over the latent state for memoryless policies. This is as if we were dealing directly with the latent MDP underlying the POMDP, a perhaps best possible scenario. Nevertheless, have we achieved exponential-free OPE in POMDPs?
@@
The answer to this question turns out to be nontrivial. In addition to the latent-state coverage parameter, the guarantee in Uehara et al. (2022a) also depends on other quantities that are less interpretable. Among them, the boundedness of FDVF itself—a concept central to this framework—is unclear, and we show that a natural construction yields an upper bound that still scales with the cumulative importance weights, thus possibly erasing the superiority of the framework over IS or MDP-reduction.
@@
    With another novel concept called belief coverage (Section 5.1), we establish fully polynomial estimation guarantee for the algorithm in Uehara et al. (2022a). The discovery of belief coverage also leads to a novel algorithm (Section 5.3) that is analogous to MIS for MDPs.
@@ 2 Preliminaries
We consider a finite-horizon POMDP $\left\langle{H,{\mathcal{S}}=\bigcup_{h=1}^{H}{\mathcal{S}}_{h},\mathcal{A},\mathcal{O}=\bigcup_{h=1}^{H}\mathcal{O}_{h},R,\mathbb{O},\mathbb{T},d_{1}}\right\rangle$, where $H$ is the horizon, ${\mathcal{S}}_{h}$ is the latent state space at step $h$ with $|{\mathcal{S}}_{h}|=S$, $\mathcal{A}$ is the action space with $|\mathcal{A}|=A$, $\mathcal{O}_{h}$ is the observation space at step $h$ with $|\mathcal{O}_{h}|=O$ , $R:\mathcal{O}\times\mathcal{A}\rightarrow[0,1]$ is the reward function, $\mathbb{O}:{\mathcal{S}}\rightarrow\Delta(\mathcal{O})$ is the emission dynamics with $\mathbb{O}(\cdot|s_{h})$ supported on $\mathcal{O}_{h}$ for $s_{h}\in{\mathcal{S}}_{h}$, $\mathbb{T}:{\mathcal{S}}\times\mathcal{A}\to\Delta({\mathcal{S}})$ is the dynamics ($\mathbb{T}(\cdot|s_{h},a_{h})$ is supported on ${\mathcal{S}}_{h+1}$), and $d_{1}\in\Delta({\mathcal{S}}_{1})$ is the initial latent state distribution. For mathematical convenience we assume all the spaces are finite and discrete, but the cardinality of $\mathcal{O}_{h}$, $O$, can be arbitrarily large. A trajectory (or episode) is sampled as $s_{1}\sim d_{1}$, then $o_{h}\sim\mathbb{O}(\cdot|s_{h})$, $r_{h}=R(o_{h},a_{h})$, $s_{h+1}\sim\mathbb{T}(\cdot|s_{h},a_{h})$ for $1\leq h\leq H$, with $a_{1:H}$ decided by the decision-making agent, and the episode terminates after $a_{H}$. $s_{1:H}$ are latent and not observable to the agent.
@@
Given an episode $o_{1},a_{1},\ldots,o_{H},a_{H}$ and a time step $h$ of interest, it will be convenient to rewrite the episode as $(\tau_{h}\,,\,\overbrace{o_{h}\,,\,a_{h}\,,\,f_{h+1}}^{f_{h}}).$ Here $\tau_{h}=(o_{1},a_{1},\ldots o_{h-1},a_{h-1})\in\mathcal{H}_{h}:=\prod_{h^{\prime}=1}^{h-1}(\mathcal{O}_{h}\times\mathcal{A})$ denotes the historical observation-action sequence (or simply history) prior to step $h$, and $f_{h+1}=(o_{h+1},a_{h+1},\ldots,o_{H},a_{H})\in\mathcal{F}_{h+1}:=\prod_{h^{\prime}=h+1}^{H}(\mathcal{O}_{h^{\prime}}\times\mathcal{A})$ denotes the future after step $h$. This format will be convenient for reasoning about the system dynamics at step $h$. We use $\mathcal{H}=\bigcup_{h=1}^{H}\mathcal{H}_{h}$ to denote the entire history domain and use $\mathcal{F}=\bigcup_{h=1}^{H}\mathcal{F}_{h}$ to denote the entire future domain. This way we can use stationary notation for functions over ${\mathcal{S}},\mathcal{O},\mathcal{H},\mathcal{F}$, where the time step can be identified from the function input (e.g., $R(o_{h},a_{h})$), and functions with time-step subscripts refer to their restriction to the $h$-th step input space, often treated as a vector (e.g., $R_{h}\in\mathbb{R}^{\mathcal{O}_{h}\times\mathcal{A}}$).
Memoryless and History-dependent Policies.
A policy $\pi:\bigcup_{h=1}^{H}(\mathcal{H}_{h}\times\mathcal{O}_{h})\rightarrow\Delta(\mathcal{A})$ specifies the action probability conditioned on the past observation-action sequence. [^1] Here we do not allow the policy to depend on the latent state, which satisfies sequential ignorability and eliminates data confoundedness; see Namkoong et al. (Section 3; 2020) and Uehara et al. (Appendix B; 2022a). There is also a line of research on OPE in confounded POMDPs where the behavior policy only depends on the latent state (Tennenholtz et al., 2020; Shi et al., 2022); see Appendix A. In the main text, we will restrict ourselves to memory-less (or reactive) policies that only depends on the current observation $o_{h}$; extension to general policies is similar to Uehara et al. (2022a) and discussed in Appendix B.6. For any $\pi$, we use $\Pr_{\pi}$ and $\mathbb{E}_{\pi}$ for the probabilities and expectations under episodes generated by $\pi$, and define $J(\pi)$ as the expected cumulative return: $J(\pi):=\mathbb{E}_{\pi}\left[\sum_{h=1}^{H}R(o_{h},a_{h})\right]$. For memoryless $\pi$, we also define $V_{{\mathcal{S}}}^{\pi}(s_{h})$ as the latent state value function at $s_{h}$: $V_{{\mathcal{S}}}^{\pi}(s_{h}):=\mathbb{E}_{\pi}\left[\sum_{h^{\prime}=h}^{H}R(o_{h^{\prime}},a_{h^{\prime}})\mid s_{h}\right]\in[0,H]$. $d^{\pi}(s_{h})$ denotes the marginal distribution of $s_{h}$ under $\pi$.
@@
In OPE, the goal is to estimate $J(\pi_{e})$ using $n$ data trajectories $\mathcal{D}=\{(o_{1}^{(i)},a_{1}^{(i)},r_{1}^{(i)},\ldots,o_{H}^{(i)},a_{H}^{(i)},r_{H}^{(i)}):i\in[n]\}$ collected using $\pi_{b}$. We write $\mathbb{E}_{\mathcal{D}}[\cdot]$ to denote empirical approximation of expectation using $\mathcal{D}$. Define the one-step action probability ratio $\mu(o_{h},a_{h}):=\frac{\pi_{e}(a_{h}\mid o_{h})}{\pi_{b}(a_{h}\mid o_{h})}$. We make the following assumption throughout:
@@
We assume $\pi_{b}(a_{h}|o_{h})$ is known and $\max_{h,o_{h},a_{h}}\mu(o_{h},a_{h})\leq C_{\mu}$.
This is a standard assumption in the OPE literature, and is needed by IS and value-based estimators that model state value functions (Jiang and Li, 2016; Liu et al., 2018).
@@
For mathematical conveniences, we make the following assumptions throughout merely for simplifying presentations; they allow us to invert certain covariance matrices and avoid $/0$ 0 situations, which can be easily handled with extra care when the assumptions do not hold.
@@
$\forall h\in[H]$, (1) $\textrm{rank}(M_{\mathcal{H},h})=\textrm{rank}(M_{\mathcal{F},h})=S$.
(2) $\forall f_{h}$,  $\Pr_{\pi_{b}}(f_{h})>0$; $\forall o_{h},a_{h}$, $R(o_{h},a_{h})>0$.
@@ 3 Future-dependent Value Functions
In this section, we provide a recap of FDVFs and translate the main result of Uehara et al. (2022a) into the finite-horizon setting, which is mathematically cleaner and more natural in many aspects; see Appendix B.2 for further discussion.
@@
The central idea is to find observable proxies of $V_{{\mathcal{S}}}^{\pi_{e}}(s_{h})$, which takes future as inputs:
Definition 3 (Future-dependent value functions (Uehara et al., 2022a)).
A future-dependent value function $V_{\mathcal{F}}:\mathcal{F}\to\mathbb{R}$, where $\mathcal{F}:=\bigcup_{h}\mathcal{F}_{h}$, is any function that satisfies the following: $\forall s_{h}$, $\mathbb{E}_{\pi_{b}}\left[V_{\mathcal{F}}(f_{h})\mid s_{h}\right]=V_{{\mathcal{S}}}^{\pi_{e}}(s_{h}).$ Equivalently, in matrix form, we have $\forall h$,
$M_{\mathcal{F},h}\times V_{\mathcal{F},h}=V_{{\mathcal{S}},h}^{\pi_{e}}.$ (1)
Recall our convention, that $V_{\mathcal{F},h}\in\mathbb{R}^{|\mathcal{F}_{h}|}$ is $V_{\mathcal{F}}$ restricted to $\mathcal{F}_{h}$, and $V_{{\mathcal{S}},h}^{\pi_{e}}$ is defined similarly.
A FDVF $V_{\mathcal{F}}$ is a property of $\pi_{e}$, but also depends on $\pi_{b}$. As we will see later in Section 4, the boundedness of $V_{\mathcal{F}}$ will depend on certain notion of coverage of $\pi_{b}$ over $\pi_{e}$. As another important property, the FDVF $V_{\mathcal{F}}$ is generally not unique even if we fix $\pi_{e}$ and $\pi_{b}$, as Eq.(1) is generally an underdetermined linear system ($S\ll|\mathcal{F}_{h}|$) and can yield many solutions. As we see below, it suffices to model any one of the solutions. Thus, from now on, when we talk about the boundedness of $V_{\mathcal{F}}$, we always consider the $V_{\mathcal{F}}$ with the smallest range among all solutions.
@@
$\forall V:\mathcal{F}\to\mathbb{R}$, the Bellman residual on state $s_{h}$ is: [^2] We let $V(f_{H+1})\equiv 0$ for all $V$ to be considered, so that we do not need to handle the $H$-th step separately.
@@
$(\mathcal{B}^{\mathcal{H}}V)(\tau_{h}):=\mathbb{E}_{\begin{subarray}{c}a_{h}\sim\pi_{e}\\ a_{h+1:H}\sim\pi_{b}\end{subarray}}[r_{h}+V(f_{h+1})\mid\tau_{h}]-\mathbb{E}_{\pi_{b}}[V(f_{h})\mid\tau_{h}]=\langle\mathbf{b}(\tau_{h}),\mathcal{B}^{{\mathcal{S}}}_{h}V\rangle.$ (2)
@@
For any $\pi_{e}$, $\pi_{b}$, and $V:\mathcal{F}\to\mathbb{R}$, $J(\pi_{e})-\mathbb{E}_{\pi_{b}}[V(f_{1})]=\textstyle\sum_{h=1}^{H}\mathbb{E}_{\pi_{e}}\left[(\mathcal{B}^{{\mathcal{S}}}V)(s_{h})\right].$
See proof in Appendix C.1. As the lemma shows, any $V$ with small $\mathcal{B}^{{\mathcal{S}}}V$ (such as $V_{\mathcal{F}}$, since $\mathcal{B}^{{\mathcal{S}}}V_{\mathcal{F}}\equiv 0$) can be used to estimate $J(\pi_{e})$ via $\mathbb{E}_{\pi_{b}}[V(f_{1})]$. But again, $\mathcal{B}^{{\mathcal{S}}}$ operates on ${\mathcal{S}}$ which is unobserved, and we turn to its proxy $\mathcal{B}^{\mathcal{H}}$, which are linear measures of $\mathcal{B}^{{\mathcal{S}}}$ (Eq.2). More concretely, Uehara et al. (2022a) proposed to estimate $\mathbb{E}_{\pi_{b}}[(\mathcal{B}^{\mathcal{H}}V)^{2}]$ using an additional helper class $\Xi:\mathcal{H}\to\mathbb{R}$ to handle the double-sampling issue (Antos et al., 2008; Dai et al., 2018):